# Conan: A Chunkwise Online Network for Zero-Shot Adaptive Voice Conversion

### v3

@@ III-E Training and Inference Procedures
Figure 3: The chunk-wise online inference procedure of Conan. Here, we illustrate the causal setting without utilizing any right context.
@@
Inference Procedure. As shown in Figure 3, we first feed the entire reference speech into the model to provide timbre and stylistic information. During chunk‑wise online inference, we wait until the input reaches a predefined chunk size before passing it to the model. Because our generation speed for each chunk is faster than the chunk’s duration, online generation becomes possible. To ensure temporal continuity, we employ a sliding context window strategy. At each generation step, we not only input the source speech of the current chunk but also include the preceding context. From the model’s output, we extract only the segment for this chunk. As the context covers the receptive field, consistent overlapping segments can be generated, ensuring smooth transitions at chunk boundaries.
@@ IV-A Experimental Setup
Evaluation Details. We objectively and subjectively evaluate our model based on three criteria: real‐time content accuracy, quality, and speaker similarity. Content accuracy is measured by Word Error Rate (WER) and Character Error Rate (CER) using a HuBERT‐Large ASR model  [^1] https://github.com/resemble-ai/Resemblyzer. Speaker similarity (SIM) is assessed with Resemblyzer  [^2] https://huggingface.co/facebook/hubert-large-ls960-ft. For subjective metrics, we conduct mean opinion score (MOS) and comparative mean opinion score (CMOS) evaluations. Both metrics are rated on a 1–5 scale and reported with 95% confidence intervals. Specifically, we use MOS‑Q to judge quality (clarity and naturalness) and MOS‑S to assess speaker similarity (timbre and styles) [47]. For CMOS‑Q and CMOS‑S, listeners compare pairs of speech samples produced by different systems. We randomly selected at least 20 pairs with more than 6 unseen speakers for each subjective evaluation, with at least 15 listeners per pair. All participants were fairly compensated and provided informed consent for the use of their responses in scientific research.
@@ IV-B Main Results
TABLE I: Evaluation results of Conan and baseline models, including both objective (WER, CER, SIM) and subjective (MOS-S, MOS-Q) evaluations. All baselines, except for StreamVC, are evaluated in an offline setting.
Method Method Content Accuracy Speaker Similarity Quality
@@
Data Source 5.41% 2.43% - - 3.96 $\pm$ 0.08
Source (Vocoder) 5.56% 2.67% - - 3.94 $\pm$ 0.15
Reference (Vocoder) - - 95.81% 4.43 $\pm$ 0.14 -
@@
As shown in Table I, Conan in the full setting delivers the highest speaker similarity (SIM, MOS-S), indicating that the Adaptive Style Encoder and timbre encoder effectively model the speaker identity and therefore achieve superior zero-shot similarity. Conan also attains the best quality (MOS-Q), demonstrating that the streaming framework enhances synthesis quality. For content accuracy, Conan records the lowest WER, while its CER is slightly higher than that of StreamVC. This is mainly because StreamVC reuses the source pitch and energy values, giving it a CER advantage. Finally, the metrics for Source (Vocoder) and Reference (Vocoder) are both excellent, further highlighting the outstanding online reconstruction capability of the Causal Shuffle Vocoder.
@@
As shown in Table II, in the faster setting, Conan uses a 20 ms chunk size to perform strictly causal, chunk‐wise online voice conversion with only 37 ms of latency, yet still achieves high performance metrics. Under Conan’s full setting with 80 ms chunk size and 2 right context chunks, we achieve high quality while still meeting real‐time latency constraints.
@@ IV-C Ablation Study
As shown in Table III, we conduct ablation studies to demonstrate the effectiveness of various design choices within the full setting. When the original HuBERT is used instead of the Stream Content Encoder, we observe a substantial deterioration in both synthesis quality and WER, underscoring the incompatibility of HuBERT with streaming scenarios. Furthermore, when right‐context chunks are omitted, all metrics worsen, indicating that the size of the lookahead can be adjusted as a trade‐off between efficiency and quality. Furthermore, removing the Adaptive Style Encoder results in a marked decrease in MOS-S, confirming that our approach effectively captures speaker identity. Finally, substituting the Causal Shuffle Vocoder with a zero-padding causal HiFi-GAN leads to a decline in MOS-Q, as the zero padding causal HiFi-GAN causes a decline in synthesis quality.
@@ V Conclusion
In this paper, we introduce Conan, a streaming voice conversion method that preserves the linguistic content of a source utterance while adopting the speaker information of a reference utterance. At its core, Conan employs the Streaming Content Extractor based on Emformer to encode source audio with minimal latency, ensuring semantic information is captured in real time. To transfer the reference speaker identity, Adaptive Style Encoder analyzes the reference speech and derives detailed speaker representations that guide the conversion process. Finally, Causal Shuffle Vocoder, built upon a fully causal HiFiGAN architecture enhanced with pixel-shuffle operations, reconstructs high-fidelity audio with low generation delay. Experimental results show that Conan surpasses existing baseline approaches, achieving better performance according to both subjective and objective metrics.

### v4

@@ III-E Training and Inference Procedures
Figure 3: The chunkwise online inference procedure of Conan. Here, we illustrate the causal setting without utilizing any right context.
@@
Inference Procedure. As shown in Figure 3, we first feed the entire reference speech into the model to provide timbre and stylistic information. During chunkwise online inference, we wait until the input reaches a predefined chunk size before passing it to the model. Because our generation speed for each chunk is faster than the chunk’s duration, online generation becomes possible. To ensure temporal continuity, we employ a sliding context window strategy. At each generation step, we not only input the source speech of the current chunk but also include the preceding context. From the model’s output, we extract only the segment for this chunk. As the context covers the receptive field, consistent overlapping segments can be generated, ensuring smooth transitions at chunk boundaries.
@@ IV-A Experimental Setup
Evaluation Details. We objectively and subjectively evaluate our model based on three criteria: real‐time content accuracy, quality, and speaker similarity. Content accuracy is measured by Word Error Rate (WER) and Character Error Rate (CER) using a HuBERT‐Large ASR model. Speaker similarity (SIM) is assessed with Resemblyzer. For subjective metrics, we conduct mean opinion score (MOS) and comparative mean opinion score (CMOS) evaluations. Both metrics are rated on a 1–5 scale and reported with 95% confidence intervals. Specifically, we use MOS‑Q to judge quality (clarity and naturalness) and MOS‑S to assess speaker similarity (timbre and styles) [45]. For CMOS‑Q and CMOS‑S, listeners compare pairs of speech samples produced by different systems. We randomly selected at least 20 pairs with more than 6 unseen speakers for each subjective evaluation, with at least 15 listeners per pair. All participants were fairly compensated and provided consent for the use of their responses in scientific research.
@@ IV-B Main Results
TABLE I: Evaluation results of Conan and baseline models, including both objective (WER, CER, SIM) and subjective (MOS-S, MOS-Q) evaluations. A tick (✓) in the Online column indicates streaming/online processing.
Category Method Online Content Accuracy Speaker Similarity Quality
@@
Data Source 5.41% 2.43% 3.96 $\pm$ 0.08
Source (Vocoder) 5.56% 2.67% 3.94 $\pm$ 0.15
Reference (Vocoder) 95.81% 4.43 $\pm$ 0.14
@@
As shown in Table I, Conan in the full setting delivers the highest speaker similarity (SIM, MOS-S), indicating that the Adaptive Style Encoder and timbre encoder effectively model the speaker timbre and therefore achieve superior zero-shot similarity. Conan also attains the best quality (MOS-Q), demonstrating that the streaming framework enhances synthesis quality. For content accuracy, Conan records the lowest WER, while its CER is slightly higher than that of StreamVC. This is mainly because StreamVC reuses the source pitch and energy values, giving it a CER advantage. Finally, the metrics for Source (Vocoder) and Reference (Vocoder) are both excellent, further highlighting the outstanding online reconstruction capability of the Causal Shuffle Vocoder.
@@
As shown in Table II, in the fast setting, Conan uses a 20 ms chunk size to perform strictly causal, chunkwise online voice conversion with only 37 ms of latency, yet still achieves high performance metrics. Under Conan’s full setting with 80 ms chunk size and 2 right context chunks, we achieve high quality while still meeting real‐time latency constraints.
@@ IV-C Ablation Study
As shown in Table III, we conduct ablation studies to demonstrate the effectiveness of various design choices within the full setting. When the original HuBERT is used instead of the Stream Content Encoder, we observe a substantial deterioration in both synthesis quality and WER, underscoring the incompatibility of HuBERT with streaming scenarios. Furthermore, when right‐context chunks are omitted, all metrics worsen, indicating that the size of the lookahead can be adjusted as a trade‐off between efficiency and quality. Furthermore, removing the Adaptive Style Encoder results in a marked decrease in MOS-S, confirming that our approach effectively captures speaker representation. Finally, substituting the Causal Shuffle Vocoder with a zero-padding causal HiFi-GAN leads to a decline in MOS-Q, as the zero padding causal HiFi-GAN causes a decline in synthesis quality.
@@ V Conclusion
In this paper, we introduce Conan, a streaming voice conversion system that preserves the linguistic content of a source utterance while adopting the speaker information of a reference utterance. We design Streaming Content Extractor, Adaptive Style Encoder, and Causal Shuffle Vocoder for better zero-shot online VC performance. Experimental results show that Conan surpasses existing baseline approaches, achieving better performance according to subjective and objective metrics.
Acknowledgment and Disclaimer
Supported by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/Interior Business Center (DOI/IBC) contract number 140D0424C0066. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon.
The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DOI/IBC, or the U.S. Government.